0.5(x+5)-x-4/3=2/5(3x+10)
Answer by mananth(16946)

0.5(x+5)-x-4/3=2/5(3x+10)
1/2 *x +5/2 -x-4/3 = 6x/5 +4
x/2 -x-6x/5 = 4-5/2 +4/3
(5x-10x-6x)/10 = (3-30+16)/12
-11x/10 = -11/12
-11x= -11*10 / 12
x= -11 * 10 / 12 * -11
x=10/12 = 5/6